Grade 12th passPhysical Chemistry

what is the kinetic energy of one mole of an ideal gas?

ishwa
7 Years agoGrade 12th pass

2 Answers

Mahima Kanawat
7 Years ago
Dear student 
 
Kinetic energy of one mole of an ideal gas is (3/2)RT.
 
WITH REGARDS 
MAHIMA 
ASKIITIANS FORUM EXPERT 
REENU SIVADARSHINI.M
7 Years ago
Average kinetic energy  of one mole of ideal gas = 3/2 [boltzmann’s constant ] [temperature]
 blotzmann’s constant =1.38*10-23
  K=3/2K T